OpenStudy (sparklestaraa):

Emily uses her college photocopier machine to photocopy at the rate of $0.08 per page. She decides to rent a photocopier machine for $90 a year. The cost of photocopying using the rented machine is $0.03 per page. Part A: Write an inequality that can be used to calculate the number of pages that Emily should photocopy in a year so that the amount she pays for the rented machine is less than the college machine. Define the variable used. ( Part B: How many pages should Emily print in a year to justify renting the photocopier? Show your work.

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (danjs):

Money to rent <money college machine

OpenStudy (danjs):

P = # pages printed Money to rent = $90 + 0.03*P

OpenStudy (danjs):

college machine money = 0.08*P

OpenStudy (danjs):

so in order that the rented machine be cheaper 90 + 0.03*P < 0.08*P

OpenStudy (danjs):

right?

OpenStudy (danjs):

0.05*P > 90 P > 1800 pages She needs to print more than 1800 pages to make the rented machine a better deal

OpenStudy (sparklestaraa):

wait do i solve the inequality and is that new problem ur making for Part B?

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (sparklestaraa):

and the rented machine is cheaper ryt?

OpenStudy (danjs):

part A is setting up the inequality Cost to rent < cost to use college machine

OpenStudy (danjs):

90 + 0.03P < 0.08P

OpenStudy (danjs):

90 plus 3 cents per page times number or pages < 8 cents per page times number of pages P

OpenStudy (danjs):

p IS THE NUMBER OF PAGES PRINTED

OpenStudy (danjs):

solving that is part B, 90 <0.05P P>1800 pages If she prints more than 1800 pages, then it is more cost effective to rent a machine

OpenStudy (sparklestaraa):

ohhhh wait it sounds stupid :( but how did u get 1800 did u multiply 90 *0.05?

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (sparklestaraa):

overall i wrote this Part A- Money to rent < Money for the college photocopier machine P= #of pages printed Money to rent= $90 +0.03*P Cost to use the college photocopier machine = 0.08*P 90+0.03*P < 0.08*P Cost to rent<Cost to use the college photocopier machine Part B- 90<0.05*P P>1800pages If Emily will print more than 1800 pages per year than it is wiser to rent the machine. is that right?

OpenStudy (danjs):

looks perfect

OpenStudy (sparklestaraa):

:D thanks more? i'll post to a different question

OpenStudy (sparklestaraa):

can u check one on here tho ? @DanJS

OpenStudy (danjs):

sure

OpenStudy (sparklestaraa):

Jose is comparing the prices of two carpet cleaning companies. Company A charges $0.40 per square foot of carpet area and an additional $22 per day for services. Company B charges 0.60 per square foot of carpet area and an additional $18 per day for services. Part A: Write equations to represent Company A's and Company B's total charges for cleaning a certain carpet area in a day. Define the variable used in the equations. Part B: Which company would charge less for cleaning 500 square feet of carpet area? Justify your answer. Part C: How much money is saved by using the services of Company A instead of Company B for cleaning 200 square feet of carpet area?

OpenStudy (sparklestaraa):

i answered Part A- Company A: y=0.40x +22 Company B: y=0.60x+18 x=the amount of square feet of carpet. Part B- Company A: y=0.40x +22 y=0.40(500)+22 y=222 Company B: y=0.60x+18 y=0.60(500)+18 y=318 Company A would charge less because $222 dollars is less than $318 dollars for company B's services. Part C- Company A= $122 Company B= $138 $138-$122= $16 $16 dollars is saved.

OpenStudy (danjs):

looks good, for the defining variables part in part a) also say that Y= the total money charged for x square feet of carpet

OpenStudy (danjs):

y(x), y is a function of x, total charged is a function of square feet of carpet

OpenStudy (sparklestaraa):

I get it :D

OpenStudy (sparklestaraa):

@DanJS thanks so much

OpenStudy (danjs):

no prob